OpenStudy (anonymous):

ILL MEDAL!!!!!!!!!! The probability of a randomly selected employee of a company being male is 60%. The probability of the employee being less than 30 years old is 70%. If the probability of the employee being less than 30 years old given that the employee is a male is 40%, what is the probability that the employee is a male, given that the employee is less than 30 years old? A. 0.34 B. 0.42 C. 0.55 D. 0.69 E. 0.71

OpenStudy (faiqraees):

P(Age<30 l Male)=(P(Age<30)∩P(Male))/P(Male) P(Age<30 l Male)= 0.4 (P(Age<30)∩P(Male))/P(Male) = 0.4 P(Age<30)∩P(Male) = 0.4 x P(Male) P(Age<30)∩P(Male) = 0.4 x 0.6 (P(Age<30)∩P(Male) = 0.24 P( Male l Age<30)= (P(Age<30)∩P(Male))/P(Age<30) P( Male l Age<30)= 0.24/0.7 P( Male l Age<30)= 12/35